# New flexible versions of extended generalized Pareto distribution for count data

### v3

@@ 3 Distributional regression framework
In distributional regression setting, we consider the parameters of distribution $Y$ vary according to the covariate $\boldsymbol{W}=(w_{1},\cdots,w_{p})^{\top}$, modeled as $Y|\boldsymbol{w}\ \sim\text{DEGPD}(\cdot,\theta(\boldsymbol{w}))$ or $Y|\boldsymbol{w}\ \sim\text{ZIDEGPD}(\cdot,\theta(\boldsymbol{w}))$, where $\theta(\boldsymbol{w})=(\theta_{1}(\boldsymbol{w}),\dots,\theta_{d}(\boldsymbol{w}))^{\top}$. This framework represents a distributional regression model (Stasinopoulos et al.,, 2018, Carrer and Gaetan,, 2022), where each parameter $\theta_{i}(\boldsymbol{w}),i=1,\cdots,d$ is linked to covariates via an additive predictor:
@@
One can estimate the equation (17) for model $M_{2}$ and $M_{3}$ by extending the code provided for $M_{1}$ in Ahmad et al., (2024). The functions developed for fitting $M_{1}$ are easy to use for $M_{2}$ and $M_{3}$ that facilitate the use of thin plate regression splines via evgam package, which is highly advantageous for modeling complex multidimensional processes. These splines are particularly suitable for spatial modeling and the representation of interaction effects by constructing tensor products. Due to space and data limitations, this paper has not implemented the distributional regression framework, but interested readers can easily apply it.
@@ 4.1 Model based simulation study
This section presents a simulation study evaluating the performance of the maximum likelihood estimator (MLE). Different parameter settings were tested to determine the MLE for the DEGPD and ZIDEGPD models. Moreover, the scale parameter $\beta$ is permanently set to one for all models. The other parameters, $\pi,\kappa$ and $\xi$, vary with the scenarios. The sample size $n=1000$ with $10^{4}$ replications was used to estimate the parameters; this sample size choice was closer to real applications.
@@ 5.1 DEGPD models
where $L$ is the maximum likelihood, $k$ represents the number of parameters in the model, and $n$ is the total number of observations in the observed set. Based on BIC given in Table 1, model $M_{2}$ shows superior performance for the AICCR dataset across all models. Despite variations in model performance, all DEGPD models exhibited satisfactory fits. Notably, DEGPD outperformed for datasets with a moderate number of zeros. In addition, the tail index estimated by all three models is similar (i.e., 0.73), while the bulk parameter $\kappa$ varies slightly. This clearly indicates that the proposed models show flexibility in the bulk part of distribution without compromising tail behavior. In contrast, existing models (see motivating examples) fail to capture extreme tail behavior when focused on the bulk and extreme tail of the distribution in the same operation. Additionally, Figure 8 (top) illustrated the Q-Q plots of the DEGPD fitted models, confirming the excellent fit of the models. A discrete Kolmogorov–Smirnov goodness-of-fit test from package dgof applied to the AICCR data set for all three models confirms an appropriate fit by its p-values (computed by Monte Carlo simulation) given in the last column of Table 1 (Arnold and Emerson,, 2011). Further, m-observation return level plots in Figure 8 (bottom) demonstrate a good fit regardless of the return period.
Table 1: Estimated parameters of DEGPD for $M_{1},M_{2}$ and $M_{3}$ fitted to insurance claims data of New York City by maximum likelihood method. The 95% bootstrap confidence bands are provided in square brackets.
@@ 5.2 ZIEGPD
Table 2: Estimated parameters of ZIDEGPD for $M_{1},M_{2}$ and $M_{3}$ to when applied doctor visits data by Maximum likelihood method. The 95% bootstrap confidence bands are provided in square brackets.
@@
These plots visually confirm that the observed quantiles closely match the expected quantiles under the fitted distributions, indicating a robust performance of the models. At the same time, this is not seen for existing models (see Figure 1). A discrete Kolmogorov–Smirnov (K-S) goodness-of-fit test was conducted on the DV dataset for all three models to substantiate the adequacy of the model further fits. The results of this test in terms of p-values (computed by Monte Carlo simulation) are reflected in the last column of Table 2 to support the appropriateness of the fitted models. High p-values suggest no significant difference between the empirical distribution of the data and the theoretical distribution specified by the ZIDEGPD models, reinforcing the validity of our proposed models. Additionally, the m-observation return level plots depicted in Figure 9 (bottom) reveal a strong fit for the ZIDEGPD models across various return periods. These plots illustrate the expected maximum values over different time intervals, indicating that the models reliably capture the extreme value behavior of the data regardless of the return period considered. Thus, the ZIDEGPD models demonstrate strong performance in estimating the proportion of zeros and provide consistent tail index estimates.
Table 3: Estimated parameters and 95% bootstrap confidence intervals of DEGPD ($M_{3}$) and DGPD when fitted to the exceedances of betting and gaming offenses above the fixed low threshold values 1 and 3.
u=1
@@
$\text{M}_{1}$ 1.09 9.61 0.11 2265.24
$[0.89,1.40]$ [5.43, 12.67] [$\times 10^{-6}$, 0.26]
$\text{M}_{2}$ 0.01 10.6 0.07 2265.75
$[2.89\times 10^{-6},2.83]$ [5.86, 11.98] [1.36×$\times 10^{-6}$, 0.29]
$\text{M}_{3}$ 1.15 9.66 0.11 2265.34
$[0.13,1.73]$ [6.62, 18.78] [1.05×$\times 10^{-6}$, 0.28]
DGPD - 10.62 0.06 2267.96
- [8.88, 12.35] [1.05×$\times 10^{-2}$, 0.18]
u=3
@@
$\text{M}_{1}$ 1.10 9.27 0.13 1895.74
$[0.88,1.52]$ [5.75, 13.02] [1.32×$\times 10^{-6}$, 0.35]
$\text{M}_{2}$ 0.01 10.48 0.08 1896.25
$[8.19\times 10^{-7},2.89]$ [5.35, 12.13] [1.58×$\times 10^{-6}$, 0.35]
$\text{M}_{3}$ 1.10 9.8 0.11 1896.12
$[0.12,1.87]$ [5.69, 18.92] [1.07×$\times 10^{-6}$, 0.35]
DGPD - 10.49 0.08 1897.63
- [8.55, 12.43] [1.06×$\times 10^{-2}$, 0.22]
@@ 5.3 DEGPD at low threshold
To evaluate the performance of the DEGPD models at low threshold exceedances, Table 3 summarizes the parameter estimates and the BIC for the DEGPD models ($M_{1}$, $M_{2}$, and $M_{3}$) and DGPD. These models are specifically fitted to the exceedances of betting and gaming offenses above the fixed low threshold values 1 and 3. These threshold values are selected corresponding to 10% and 20% quantile of the data. According to the BIC, all DEGPD models perform better than the DGPD, which is expected due to the use of the threshold being too low for the DGPD model to be appropriate. With the DGPD model, the tail index estimate is underestimated at a lower level $\hat{\xi}=0.06(1.05\times 10^{-2},0.18)$ when the chosen threshold is very low, i.e., 1. It is consistent with the simulations (see Figure 6), indicating that the tail index of DGPD is underestimated when the threshold is too low. For DEGPD ($M_{1}$) and DEGPD ($M_{3}$), the tail indexes are less underestimated, but for DEGPD ($M_{2}$), the tail index estimation is almost the same as DGPD when the chosen threshold is 3, a bit higher than low. The BIC in Table 3 favors the DEGPD($M_{1}$) model. To observe what happens with high threshold selection, we fitted all models reported in Table 3 with a threshold equal to 7. The results are not reported here. From non-reported results, we observe that the tail indexes for DEGPD($M_{1}$) and DEGPD($M_{3}$) were underestimated, DGPD was performing better for exceedances over a high threshold, and the tail parameter of DEGPD($M_{2}$) stayed very near to DGPD tail index. Figure 10 shows the Q-Q and m-observation return level plots for DEGPD($M_{1}$) when fitted model at low threshold values 1 and 3. The plots also confirm that the model is a good fit.
@@ 6 Discussion and conclusion
This paper proposes flexible versions of DEGPD, especially suitable for dealing with the entire range of count data, a large number of zeros, and conditional exceedances over a low threshold. We presented two new families of distributions that show flexibility in the above-discussed cases: one based on the truncated Normal distribution, another based on truncated Beta distribution, and an existing one based on power law that has already been discussed in Ahmad et al., (2024). Consider distributions extending the DEGPD and ZIDEGPD models introduced in Ahmad et al., (2024) by adding new extended models for which the distributions are flexible with nonnegative integers and zero excessiveness.
@@
Furthermore, the DEGPD models could be used to develop methods for threshold selection in integer-valued data. In contrast to the traditional threshold-relying model discussed in Hitz et al., (2024), where selecting a sufficiently high threshold is crucial for the asymptotic approximation to be valid, extended models offer greater flexibility near the threshold. Therefore, threshold selection is less critical in extended models, as their probability mass functions can adapt more readily to variations near the threshold.
@@
Acknowledgement(s)
Touqeer Ahmad acknowledges support from the Règion Bretagne through project SAD-2021- MaEVa.

### v4

@@ 3 Distributional regression framework
In distributional regression setting, we consider the parameters of distribution $Y$ vary according to the covariate $\boldsymbol{W}=(w_{1},\cdots,w_{p})^{\top}$, modeled as $Y|\boldsymbol{w}\ \sim\text{DEGPD}(\cdot,\theta(\boldsymbol{w}))$ or $Y|\boldsymbol{w}\ \sim\text{ZIDEGPD}(\cdot,\theta(\boldsymbol{w}))$, where $\theta(\boldsymbol{w})=(\theta_{1}(\boldsymbol{w}),\dots,\theta_{d}(\boldsymbol{w}))^{\top}$. This framework represents a distributional regression model [22, 6], where each parameter $\theta_{i}(\boldsymbol{w}),i=1,\cdots,d$ is linked to covariates via an additive predictor:
@@
One can estimate the equation (17) for model $M_{2}$ and $M_{3}$ by extending the code provided for $M_{1}$ in Ahmad et al. [1]. The functions developed for fitting $M_{1}$ are easy to use for $M_{2}$ and $M_{3}$ that facilitate the use of thin plate regression splines via evgam package, which is highly advantageous for modeling complex multidimensional processes. These splines are particularly suitable for spatial modeling and the representation of interaction effects by constructing tensor products. Due to space and data limitations, this paper has not implemented the distributional regression framework, but interested readers can easily apply it.
@@ 4.1 Model based simulation study
This section presents a simulation study evaluating the performance of the maximum likelihood estimator (MLE). Different parameter settings were tested to determine the MLE for the DEGPD and ZIDEGPD models. Moreover, the scale parameter $\beta$ is permanently set to one for all models. The other parameters, $\pi,\kappa$ and $\xi$, vary with the scenarios. A sample size of $n=1000$ with $10^{4}$ replications was used to estimate the parameters; this sample size choice was more representative of real-world applications. During the analysis, the initial values for the DEGPD parameters $c(\kappa,\beta,\xi)$ were set as $c\Big(\kappa=\sqrt[4]{\mathrm{mean}(Y)},\;\beta=\sqrt[4]{\mathrm{sd}(Y)},\;\xi=0.2\Big),$ and for the ZIDEGPD parameters $c(\pi,\kappa,\beta,\xi)$ as $c\Big(\pi=0.2,\;\kappa=\sqrt[4]{\mathrm{mean}(Y)},\;\beta=\sqrt[4]{\mathrm{sd}(Y)},\;\xi=0.2\Big),$ where $Y$ denotes the data generated from the model under consideration.
@@ 5.1 DEGPD models
where $L$ is the maximum likelihood, $k$ represents the number of parameters in the model, and $n$ is the total number of observations in the observed set. Based on the BIC given in Table 1, model $M_{2}$ shows superior performance for the AICCR dataset across all models, but higher variability is observed in parameter $\kappa$ when compared $M_{1}$ and $M_{3}$. Despite variations in model performance, all DEGPD models exhibited satisfactory fits. Notably, DEGPD outperformed for datasets with a moderate number of zeros. In addition, the tail index estimated by all three models is similar (i.e., 0.73), while the bulk parameter $\kappa$ varies slightly. This clearly indicates that the proposed models show flexibility in the bulk part of distribution without compromising tail behavior. In contrast, existing models (see motivating examples) fail to capture extreme tail behavior when focused on the bulk and extreme tail of the distribution in the same operation. Additionally, Figure 8 (top) illustrated the Q-Q plots of the DEGPD fitted models, confirming the excellent fit of the models. A discrete Kolmogorov–Smirnov goodness-of-fit test from package dgof applied to the AICCR data set for all three models confirms an appropriate fit by its p-values (computed by Monte Carlo simulation) given in the last column of Table 1 [2]. Further, m-observation return level plots in Figure 8 (bottom) demonstrate a good fit regardless of the return period.
Table 1: Estimated parameters of DEGPD for $M_{1},M_{2}$ and $M_{3}$ fitted to insurance claims data of New York City by maximum likelihood method. The 95% bootstrap confidence bands are provided in square brackets, and standard errors are given in parentheses.
@@ 5.2 ZIEGPD
Table 2: Estimated parameters of ZIDEGPD for $M_{1},M_{2}$ and $M_{3}$ when applied to doctor visits data by the Maximum likelihood method. The 95% bootstrap confidence bands are provided in square brackets, and standard errors are given in parentheses.
@@
These plots visually confirm that the observed quantiles closely match the expected quantiles under the fitted distributions, indicating a robust performance of the models. At the same time, this is not seen for existing models (see Figure 1). A discrete Kolmogorov–Smirnov (K-S) goodness-of-fit test was conducted on the DV dataset for all three models to substantiate the adequacy of the model’s further fits. The results of this test in terms of $p$-values (computed by Monte Carlo simulation) are reflected in the last column of Table 2 to support the appropriateness of the fitted models. High $p$-values suggest that there is no significant difference between the empirical distribution of the data and the theoretical distribution specified by the ZIDEGPD models, thereby reinforcing the validity of our proposed models. Additionally, the m-observation return level plots depicted in Figure 9 (bottom) reveal a strong fit for the ZIDEGPD models across various return periods. These plots illustrate the expected maximum values over different time intervals, indicating that the models reliably capture the extreme value behavior of the data regardless of the return period considered. Thus, the ZIDEGPD models demonstrate strong performance in estimating the proportion of zeros and provide consistent tail index estimates.
Table 3: Estimated parameters and 95% bootstrap confidence intervals of DEGPD ($M_{3}$) and DGPD when fitted to the exceedances of betting and gaming offenses above the fixed low threshold values 1 and 3. Bootstrap confidence bands are provided in square brackets, and standard errors are given in parentheses.
u=1{u=1}
@@
$\text{M}_{1}$ 1.09 (0.14) 9.61 (1.58) 0.11 (0.07) 2265.24
[…]
$\text{M}_{2}$ 0.01 (0.75) 10.6 (1.51) 0.07 (0.08) 2265.75
[…]
$\text{M}_{3}$ 1.15 (0.35) 9.66 (2.69) 0.11 (0.08) 2265.34
[…]
DGPD - 10.62 (1.53) 0.06 (0.08) 2267.96
[…]
u=3{u=3}
@@
$\text{M}_{1}$ 1.10 (0.17) 9.27 (2.00) 0.13 (0.10) 1895.74
[…]
$\text{M}_{2}$ 0.01 (0.73) 10.48 (1.56) 0.08 (0.09) 1896.25
[…]
$\text{M}_{3}$ 1.10 (0.40) 9.80 (3.10) 0.11 (0.10) 1896.12
[…]
DGPD - 10.49 (1.57) 0.08 (0.10) 1897.63
[…]
@@ 5.3 DEGPD at low threshold
To evaluate the performance of the DEGPD models at low threshold exceedances, Table 3 summarizes the parameter estimates and the BIC for the DEGPD models ($M_{1}$, $M_{2}$, and $M_{3}$) and DGPD. These models are specifically fitted to the exceedances of betting and gaming offenses above the fixed low threshold values 1 and 3. These threshold values are selected corresponding to the 10% and 20% quantiles of the data. According to the BIC, all DEGPD models outperform the DGPD, which is expected due to the threshold being too low for the DGPD model to be appropriate. With the DGPD model, the tail index estimate is underestimated at a lower level $\hat{\xi}=0.06(1.05\times 10^{-2},0.18)$ when the chosen threshold is very low, i.e., 1. It is consistent with the simulations (see Figure 6), indicating that the tail index of DGPD is underestimated when the threshold is too low. For DEGPD ($M_{1}$) and DEGPD ($M_{3}$), the tail indexes are less underestimated, but for DEGPD ($M_{2}$), the tail index estimation is almost the same as DGPD when the chosen threshold is 3, a bit higher than low. The BIC in Table 3 favors the DEGPD($M_{1}$) model. To observe what happens with high threshold selection, we fitted all models reported in Table 3 with a threshold equal to 7. The results are not reported here. From non-reported results, we observe that the tail indexes for DEGPD($M_{1}$) and DEGPD($M_{3}$) were underestimated, DGPD was performing better for exceedances over a high threshold, and the tail parameter of DEGPD($M_{2}$) stayed very near to DGPD tail index. Figure 10 shows the Q-Q and m-observation return level plots for DEGPD($M_{1}$) when fitted model at low threshold values 1 and 3. The plots also confirm that the model is a good fit.
@@ 6 Discussion and conclusion
This paper proposes flexible versions of DEGPD, especially suitable for dealing with the entire range of count data, a large number of zeros, and conditional exceedances over a low threshold. We presented two new families of distributions that show flexibility in the above-discussed cases: one based on the truncated Normal distribution, another based on truncated Beta distribution, and an existing one based on power law that has already been discussed in Ahmad et al. [1]. Consider distributions extending the DEGPD and ZIDEGPD models introduced in Ahmad et al. [1] by adding new extended models for which the distributions are flexible with nonnegative integers and zero excessiveness.
@@
Furthermore, the DEGPD models could be used to develop methods for threshold selection in integer-valued data. In contrast to the traditional threshold-relying model discussed in Hitz et al. [12], where selecting a sufficiently high threshold is crucial for the asymptotic approximation to be valid, extended models offer greater flexibility near the threshold. Therefore, threshold selection is less critical in extended models, as their probability mass functions can adapt more readily to variations near the threshold.